lance armstrong won his fourth straight tour de France bike race in 2002. he is able to travel 32 miles per hour without resting. at this rate, how many feet per second he would be traveling?
Found 2 solutions by solver91311, Alan3354:
Answer by solver91311(12126)



There are 5280 feet in a statute mile.

There are 60 minutes in an hour and 60 seconds in a minute so there are



seconds in an hour.

So multiply 32 miles per hour times:




John


Answer by Alan3354(21605)

lance armstrong won his fourth straight tour de France bike race in 2002. he is able to travel 32 miles per hour without resting. at this rate, how many feet per second he would be traveling?
---------------
15 mph = 22 ft/sec
1 mph = 22/15 fps
32 mph = 32(22/15) fps
32 mph = 704/15 fps
32 mph = 46.93333... fps